OpenStudy (anonymous):

there is an 8 percent chance it will rain tomorrow. what is the probability that it wont rain tomorrow?

OpenStudy (anonymous):

percentage is out of 100, so to get the chance use the maximum percentage minus the chance (100 - 8)

OpenStudy (anonymous):

^ What she said. You can also think of it as fractions. 1 = (chance it will rain tomorrow) + (chance it won't) 1 = (0.08) + x x = 0.92 So there's a 92 percent chance of it not raining. Remember that total probability is 100% or 1